1.       What volume of 0,250 mol⋅dm–3 H2SO4 (sulfuric acid) will react with 42,4 cm3 of 1,50 mol⋅dm-3 Na2CO3 (sodium carbonate)? (Hint: Write the balanced equation first.) 


1
Expert's answer
2021-04-22T03:19:06-0400

H2SO4 + Na2CO3 =H2O + CO2 + Na2SO4


n(Na2CO3)=0.042×1.5=0.06mol

n(H2SO4)=n(Na2CO3)


V(Na2CO3)=0.06/0.250= 0.25l=250ml